Status Report

ISS On-Orbit Status 28 Jan 2003

By SpaceRef Editor
January 28, 2003
Filed under , ,
ISS On-Orbit Status 28 Jan 2003

All ISS systems continue to function nominally, except as noted previously
or below. Today 17 years ago, we lost the Shuttle "Challenger" with
its crew of seven: Dick Scobee, Michael Smith, Ronald McNair, Ellison
Onizuka, Judith Resnick, Greg Jarvis and Christa McAuliffe. Also sadly
remembered: the loss of Roger Chaffee, Ed White and Gus Grissom in the
Apollo 1 fire, yesterday 36 years ago.

Onboard the ISS, the crew began trash stowage on 9P/Progress M1-258,
accompanied by IMS (inventory management system) audit and video recording
of the activities with the U.S. Sony PD 100 DVCAM. The video recording
was downlinked at 12:20pm EST. Deorbit of the cargo ship is still set
for next Saturday, 2/1 (separation 10:59am EST, entry interface 2:47pm).

The crew completed the scheduled fit check of the "Kazbeks",
the three contoured shock-absorbing seats in the Soyuz descent capsule
(SA). This required them to don their Sokol pressure suits, take their
seats and measure the gap between the top of the head and the top edge
of the structure facing the head with a ruler. The results were reported
to MCC-M. [The Kazbek-Us are designed to withstand g-loads during launch
and orbital insertion as well as during reentry and brake-rocket-assisted
landing. Each seat has two positions: cocked (armed) and noncocked. In
the cocked position, they are raised to allow the shock absorbers to
function during touchdown. The fit check assures that the crew, whose
bodies gain in length during longer-term stay in zero-G, would still
be adequately protected by the seat liners in case of an emergency return.]

In the DC-1 "Pirs" docking module, FE-1 Nikolai Budarin performed
the regular (monthly) check-up on the circuit breakers (AZS) on the BVP
amp switch panel — they should all be On — and the 14 LEDs of the fuses
in fuse panels BPP-30 and BPP-36.

FE-2/SO Don Pettit worked on EXPRESS rack 3, preparing its ARIS (active
rack isolation system) for testing with the new software load, adjusting
the snubbers (pin/cup) to prevent hyperextension of the shock absorber
pushrods, then calibrating the pushrods.

This was the last day for the current round of the Renal Stone experiment
for Expedition 6, with Nikolai Budarin terminating his sample collections
after breakfast and later stowing the equipment. [The Renal research,
which investigates how to prevent kidney stone formation in zero-G, is
a NASA/JSC program with Dr. Peggy Whitson as Principal Investigator (PI).]

Don Pettit reconfigured power and data cables for the EarthKAM payload,
then unstowed, set up and activated the EarthKAM system. [EarthKAM (Earth
Knowledge Acquired by Middle school students)) is using an ESC 460C electronic
still camera mounted at the Lab’s down-facing window, equipped with a
50mm lens (f1.4). Powered by 16Vdc from a 120 Vdc adapter, EarthKAM takes
pictures by remote operation from the ground, without crew interaction.
After its activation and checkout by the ground, it will be available
for Middle School-ers, who will submit image requests and conduct geographic
research. Currently, there are 27 middle schools slated to submit requests,
including 10 schools new to the program. The requests will be uplinked
in a camera control file to an SSC (station support computer) laptop
which then activates the camera at specified times and receives the digital
images from the camera’s storage card on its hard drive, for subsequent
downlink via OPS LAN. Images are then immediately posted on the Internet
at which already displays stunning imagery from previous EarthKAM flights on ISS.
The current EarthKAM run is the last as a small pressurized payload; next time
EarthKAM will operate from the WORF (window observational research facility).
Principal investigator is Dr. Sally Ride, UCSD (University of California, San

Budarin completed his regular daily inspection of the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

CDR Ken Bowersox performed the daily routine tasks of SOSh life support
systems maintenance and IMS delta file preparation for downlink, while
Pettit took the periodic CO2 (carbon dioxide) readings of the ACS (atmosphere
control and supply) in SM and Lab with the portable CDM (CO2 monitor).

At 10:05am EST, MCC-M/TsUP began with the scheduled vacuum purge of
the Progress 9P fuel (ZUG) and oxidizer (ZUO) lines, to vent prop residuals
in the plumbing between Progress and SM into space. [Fuel purge (unsymmetrical
dimethyl hydrazine, UDMH) began at 10:10am, oxidizer purge (nitrogen
tetroxide, N2O4) at 11:40am, for about 13 minutes. Planned video recording
of the oxidizer line venting was cancelled.]

Later in the day, Nikolai Budarin set up the Russian MO-21 "Ecosphera" air
sampler and incubation equipment, for atmospheric microbial air sampling
scheduled for tomorrow, and started recharge of its power supply unit.
[MO-21 determines microbial contamination of the ISS atmosphere, specifically
the total bacterial and fungal microflora counts and microflora composition
according to morphologic criteria of microorganism colonies.]

All crewmembers performed their daily physical exercise, and Bowersox
completed the weekly checkup of the TVIS treadmill (a time and date check
done just prior to powerdown or end-of-session).

MCC-M/TsUP was scheduled to conduct a data transmission test to the
onboard central computer system via the FGB "Komparus" command
and measuring system (KIS).

The crew was complimented on the Node starboard hatch window installation
which they completed successfully over the weekend. [The leak rate was
calculated to be less than six thousandth cubic centimeter per second,
well within leak rate specifications. The go-ahead was given to tear
down the VAJ (vacuum access jumper) leak check equipment and to close
up the porthole frames with a blind gland.]

A 72-hour leak check on the high-pressure side of the newly (1/18) replaced
O2 regulator in the Airlock to the PCA (pressure control assembly) was
successfully concluded today. The regulator is sealing correctly.

The S1 radiator clearance check on 1/27, which included TRRJ (thermal
radiator rotary joint) rotation, observed with the SSRMS video camera,
was completed, showing "great clearance": an ample distance
of 14 inches between the deployed S1 radiator panel and the P6 starboard
radiator at closest approach.

Temperatures of the ETCS NTA (external thermal control system/nitrogen
tank assembly) on the P1 truss have increased satisfactorily, leveling
off at about 0 degC. [Checkout of the four NTA valves was successful;
they all opened and are working correctly. Thus, the failure of the two
Loop B heaters is a heater issue, not one of the entire assembly. Analysis
continues on all expected flight attitudes between now and 12A.1, to
ensure safe hardware temperatures.]

An RPCM (remote power controller module) in the Lab, controlling Airlock
equipment, has tripped. Preliminary data indicate no overcurrent event,
which could have blown the fuse, but further analysis is needed to determine
whether it is an RPC issue or a downstream-equipment issue. The RPCM
is expected to be reset soon.

Avionics specialists have delivered all software products for the upcoming
command and control software upgrade. The new software will be available
on orbit in May.

U.S. and Russian contamination experts have finally agreed on a joint
water venting protocol, now ready for signing by both sides. [It allows
the U.S. to use a six-day window after Progress 10P docking for nonpropulsive
venting, without expending any maneuver propellants. No venting will
be done when a Russian docking port is open. We also can vent nonpropulsively
from +/-YVV/"Barbecue" attitude, when required at solar Beta
angles larger than 60 degrees. Discussions are also getting underway
between both sides on increasing condensate water collection on the Russian
segment and its proper certification. Pending necessary improvements
to the Russian SRV-K water processor system, after 10P, U.S. water processing
by the Russian Elektron should also be resumed.]

Yesterday’s SM/Kurs rendezvous test for the Progress 10P docking on
2/4 was successful. The test involved a new procedure, which uses only
one (#1) of the two Kurs antennas mounted at the tips of the two Service
Module (SM) solar wings. [The second antenna, on wing #2, is operational
but functions only intermittently. Antenna selection is done automatically
by the Kurs avionics, and yesterday’s test showed that the docking can
be performed with only the #1 antenna, made possible by the current LVLH
flight attitude which has the #1 solar wing always pointed to the approaching
vehicle. In case of Kurs failure, the old manual TORU backup mode continues
to be available.]

Launch of Progress 10P (M-247) is still set for Sunday, 2/2 (7:59am
EST), with docking on 2/4 (9:48am). The two parts for the MSG (microgravity
science glovebox) are on board the resupply ship.

Today’s targets for the CEO (crew earth observations program) were Patagonian
Glaciers (glaciers that lie on the west side of the southern Andes are
less well photographed due to greater cloudiness. Trying for Pacific-side
glacier tongues in cloud holes under present relatively clear conditions.
The handheld imagery on these two ice fields is now the best archive
on the region that exists in terms of coverage and detail), Saharan dust
outblow (Dynamic Event. Major dust pall is hundreds of miles wide, streaming
west into the Atlantic. Looking right for the southern dust margin, and
then ~2-3 min later looking left for the northern margin near the Moroccan
coast. Wide views [mapping series] of the full extent of the phenomenon
are requested), La Paz (looking right on the major western lip of the
high desert plateau for this large city, which lies both on the high
plateau and cascades down into a major canyon on the Andean flank), and
Biomass burning, Australia (Dynamic Event. Look right to document the
large smoke pall that has developed from the major fires in SW Australia.
Obliques and slight overexposure help to capture the more subtle aerosol
loadings (underexposure captures the denser, core areas of the dust mass).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber
is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control
subsystem) is operating. MCA (major constituents analyzer) is operating.
BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2
in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.

SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C)
— 24.8; ppO2 (mmHg) — 169.3; ppCO2 (mmHg) — 2.5 (suspect).
SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C)
— 19.8.
FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 21.7.
Node: Pressure (mmHg) — 745.8; temperature (deg C) — 22.9 (shell);
ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab: Pressure (mmHg) — 748.05; temperature (deg C) — 22.6; ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 748.05; temperature (deg
C) — 22.0; shell heater temp (deg C) — 22.1, ppO2 (mmHg) — 165.4;
ppCO2 (mmHg) — 5.9.
PMA-1: Shell heater temp (deg C) — 23.1
PMA-2: Shell heater temp (deg C) — 17.7
(n/a = data not available)

Propulsion System (PS): Total propellant load available (SM + FGB +
Progress) — 3669 kg (8089 lb) [as of 1/23/03]. (Capability: SM — 870
kg; FGB — 6160 kg).

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B both in "blind" dual-angle mode (directed position).
SM batteries: Battery #7 is off line (failed); battery #2 is offline;
all other batteries (6) are in "Full Charge" mode.
FGB batteries (3am): Battery #2 is offline; all other batteries (5) are
in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-1 is On (primary), EXT-2 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane
1 is down (as of 11/14/02).
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line (CMG-1 failed).
State vector — U.S. SIGI-1
Attitude — U.S. SIGI-1
Angular rates — U.S. RGA-1

Flight Attitude:
LVLH TEA (local vertical/local horizontal = "earth-fixed":
z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch:
-9.1 deg, roll: 0 deg]), CMG/Thruster Assist Momentum Management).
Solar Beta Angle: 22.6 deg (magnitude decreasing).

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both
MBS: KA power on both strings.
MT: latched at WS4, with KA power.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 9:42am EST [= epoch]):
Mean altitude — 387.8 km
Apogee — 392.5 km
Perigee — 383.2 km
Period — 92.31 min.
Inclination (to Equator) — 51.63 deg
Eccentricity — 0.0006821
Orbits per 24-hr. day — 15.60
Altitude loss — 250 m (mean) in last 24 hours
Solar Beta Angle — 22.6 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 23922

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.